# ASGO: Adaptive Structured Gradient Optimization

### v1

@@ 5 Smooth Theory
It is also important to study the performance of the algorithm in smooth settings, as many real training tasks have been widely observed to be smooth, at least locally. Also, only in smooth settings can we properly describe the importance of batch size.
@@
If we denote $X=[x_{1},\dots x_{n}]$, where each $x_{i}\in\mathbb{R}^{m}$, and vectorize $X$, we obtain that
${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\begin{bmatrix}x_{1}\\ \cdots\\ x_{n}\end{bmatrix}^{\top}\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\begin{bmatrix}x_{1}\\ \cdots\\ x_{n}\end{bmatrix},$
which means that Assumption 2 is equivalent to the existence of a symmetric matrix $L\in\mathbb{R}^{m\times m},L\succ 0$ such that for any $w\in\mathbb{R}^{mn}$,
$-\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\preceq\nabla^{2}f_{v}(w)\preceq\begin{bmatrix}L&&&\\ &L&&\\ &&\cdots&\\ &&&L\end{bmatrix}\in\mathbb{R}^{mn\times mn},$
where $f_{v}(w)=f(W)$ if $W=[w_{1},\dots,w_{n}]\in\mathbb{R}^{m\times n}$ and $w=[w_{1};w_{2};\dots;w_{n}]\in\mathbb{R}^{mn}$ is the vectorization of $W$. Therefore, this assumption is closely related to the blockwise diagonal structure of the Hessian as discovered in practice (Collobert, 2004; Zhang et al., 2024a). Also note that Assumption 2 implies the standard smoothness assumption with respect to the Frobenius norm by $\left\|\nabla f_{v}(w)\right\|_{\mathrm{op}}\leq\left\|L\right\|_{\mathrm{op}}$. This block-wise diagonal smoothness assumption is a different direction for extending the smoothness condition from the diagonal anisotropic smoothness assumption used for sign-based and adaptive gradient methods (Bernstein et al., 2018; Liu et al., 2024).
@@
Let $N\triangleq\nabla f(W;\xi)-\nabla f(W_{t})\in\mathbb{R}^{m\times n}$ be the stochastic gradient noise. We assume that $\mathbb{E}[N]=0$ and there exists a symmetric positive definite matrix $V$ such that
$\mathbb{E}\left[NN^{\top}\right]\preceq V^{2}.$
One can check that Assumption 3 implies the standard variance bound by $\mathbb{E}[\left\|N_{t}\right\|_{\mathrm{F}}^{2}]\leq\left\|V\right\|_{\mathrm{F}}^{2}$. The assumption shares some similarity with the coordinate-wise variance bound as in Bernstein et al. (2018); Crawshaw et al. (2022); Liu et al. (2024), in the sense that it allows a more fine-grained analysis. This matrix-form variance upper bound may better describe the real case since it takes the structure of the noise into account, which is relevant to the rank and other structured properties.
(a) Hessian of a MLP (Collobert, 2004)
(b) Hessian of a MLP at initialization
(c) Hessian of a MLP at 50% step
(d) Hessian of a MLP at 100% step
Figure 1: This figure is from Zhang et al. (2024a, b), describing the near-block-wise diagonal Hessian structure of neural nets. (a) is the Hessian of an MLP after 1 training step reported in (Collobert, 2004). (b,c,d): the Hessians of a 1-hidden-layer MLP with 8 neurons on CIFAR-100. The near-block-diagonal structure maintains throughout training, where each block corresponds to one neuron.
@@
Under Assumptions 1, 2 and 3, for Algorithm 1 with $\eta_{t}\equiv\eta=D_{\mathrm{op}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})-f(W_{*})]\leq\frac{4D_{\mathrm{op}}^{2}\left\|L\right\|_{*}}{T}+\frac{2\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{MT}}+\frac{2\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T},$
where $M$ is the batch size, $D_{\mathrm{op}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{op}}$, and $D_{\mathrm{F}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{F}}$..
As also discussed in Section 4, Theorem 3 depends on $D_{\mathrm{op}}$ and $D_{\mathrm{F}}$, which can be bounded if we additionally employ a projection step when doing the update. Theorem 3 indicates a similar general convergence rate to Corollary 2 for a small batch size, where the $\mathcal{O}(1/\sqrt{MT})$ term should be dominant in the rate, and thus shares the same properties as we discussed in Section 4. This means that ASGO can benefit if $V$, i.e., the low-rank stochastic gradient noise, is generally low-rank. Furthermore, when the batch size $M$ is large such that the $\mathcal{O}(\left\|L\right\|_{*}/T)$ term is also significant in the bound, Theorem 3 implies more, which can be reflected in the following discussions.
  • Comparison with SGD. SGD has the convergence rate of (Garrigos and Gower, 2023):
@@
    where $D_{\mathrm{F}}$ indicates the Frobenius norm upper bound for weights. Since the comparison between the $\mathcal{O}(1/\sqrt{MT})$ term is generally consistent with the discussion in Section 4, we further compare the $\mathcal{O}(1/T)$ term here to provide some more intuitions:
    • we have $\left\|L\right\|_{\mathrm{op}}\leq\left\|L\right\|_{*}\leq r_{L}\left\|L\right\|_{\mathrm{op}}$, where $r_{L}$ is the rank of $L$. Thus when $L$ is low-rank, or have very imbalanced singular values, $\left\|L\right\|_{*}$ can be close to $\left\|L\right\|_{\mathrm{op}}$;
    • we have $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}$ is the max rank of $W_{t}-W_{*}$. Thus when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values in a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
    Therefore, we can see in general, ASGO should work well when the Hessian can be well approximated by a block-wise diagonal matrix with a low-rank $L$ for each block and $W_{t}-W_{*}$ are relatively high-rank. Note that Hessians are discovered to be low-rank in DNNs, especially after some steps of training (Sagun et al., 2016, 2017; Wu et al., 2020).
  • Block-wise diagonal structure of Neural Networks. It has been widely observed that the Hessian of MLPs, as well as deep neural networks, are approximately block-wise diagonal (Collobert, 2004; Zhang et al., 2024a, b), as shown in Figure 1. This block-wise diagonal structure naturally arises from the structure of MLPs, i.e., the use of neurons. Inside each neuron, the parameters are more closely related and thus have a denser Hessian matrix, while outside the neuron, the Hessian can be very sparse. Based on Theorem 3, our algorithm should perform well in this block Hessian setting, showing great potential in real applications.
  • Intuitions on single-side preconditioner. Since we only use a single-side preconditioner for ASGO, it is also important to determine on which side we shall apply the preconditioner. As demonstrated in Figure 1, a block in the Hessian commonly corresponds to a neuron in the next layer, which is presented as a row of the weight matrix in MLPs. Therefore, we may desire the single-side preconditioner to be on the right side of $G_{t}$ to correctly fit this block-wise diagonal property. Based on this block-wise diagonal structure, if Hessian is approximately block-wise diagonal, Shampoo-like double-side preconditioners may not only suffer from higher memory and computation costs, but also a less effective preconditioning since they are trying to approximate the structured curvature information matrix using a dense matrix produced by Kronecker product (Gupta et al., 2018; Morwani et al., 2024).
@@
The proof of Theorem 3 is available in Appendix D. The analysis follows a similar mainline as the smooth analysis for AdaGrad (Levy et al., 2018; Liu et al., 2024), but is more complex because of the involvement of matrix operations in ASGO.
6 More Discussions on the Algorithm
@@
Muon (Jordan et al., 2024) can be interpreted as a standard steepest descent algorithm utilizing the spectral norm (Bernstein and Newhouse, 2024) with momentum if we ignore the incorporated momentum, i.e.
@@
which preserves the matrix structure of the problem and should naturally exploit the structured properties because of the involvement of spectral norm in the steepest descent framework. This is consistent with what we discussed for ASGO. Moreover, if we ignore the gradient accumulation in the preconditioner and momentum, we can see that ASGO is equivalent to Muon. [^1] A proof can be found in Appendix E. Interestingly, this equivalence between ASGO and Muon in both theoretical intuition and algorithm design is analogous to that between diagonal AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). In this sense, we may interpret Muon and ASGO as SignSGD and AdaGrad for structured gradients, respectively. Note that Shampoo also admits such a relation with Muon in algorithmic design. However, as discussed in Section 4, Shampoo achieves a much worse convergence rate and thus fails to benefit from the structured properties as ASGO does. To some extent, this means that ASGO may be a more appropriate approach than Shampoo as such an analog to diagonal AdaGrad.
From this intuition, we may also conjecture the nonconvex convergence rate of ASGO. We prove the following non-convex convergence for Muon:
$\min_{0\leq t\leq T-1}\left\|\nabla f(W_{t})\right\|_{*}^{2}\leq\mathcal{O}\left(\frac{\left\|L\right\|_{*}(f(W_{0})-f^{*})}{T}\right)$ (3)
in Appendix E, where we assume that $f^{*}\triangleq\inf f(W)>-\infty$. Since diagonal AdaGrad and SignSGD achieve just the same convergence rate in nonconvex settings up to logarithmic factors (Bernstein et al., 2018; Sun et al., 2023; Liu et al., 2024), we may expect that ASGO could achieve a comparable convergence rate with (3) obtained by Muon. It is an interesting future topic to prove this conjecture and further explore the nonconvex behavior of ASGO theoretically.
@@
Note that a nonconvex smooth convergence analysis of Muon has been presented in Li and Hong (2025). They establish the convergence results under a more general setting, involving gradient noise and momentum following the analysis in Cutkosky and Mehta (2020). However, if we only look at the deterministic case, their result is worse than (3) because of the explicit dependence on dimension $n$. The key here is that the standard smoothness condition with respect to the Frobenius norm is not a good fit for analyzing structured gradient algorithms like Muon or ASGO. Using Assumption 2, we obtain better convergence results for Muon in (3).
A practical implementation of ASGO.
We present a practical implementation of ASGO in Algorithm 2. Our implementation incorporates several common modifications to enhance computational efficiency and stability. (i) Position the preconditioner on the right side, which better aligns with the neuron architecture in DNNs as discussed in Section 5.222The convergence guarantees remain valid since this modification is equivalent to applying the original ASGO with a transposition of the parameter matrix $W$. Notably, right-side preconditioning requires significantly less computational effort than left-side preconditioning, particularly for 1D parameters such as bias terms and layer normalization weights. Left-side preconditioning would correspond to full-matrix AdaGrad in the 1D case, resulting in higher memory consumption and computational complexity. In contrast, right-side preconditioning only requires maintaining a scalar value and implementing AdaGrad-Norm (Streeter and McMahan, 2010; Ward et al., 2020), a scalar step size version of AdaGrad. This approach enables ASGO to efficiently handle all parameter structures within a unified framework, unlike methods such as Muon that require defining separate optimizers (e.g., AdamW (Loshchilov and Hutter, 2017)) for 1D parameters. Although we have not empirically verified whether this AdaGrad-Norm update rule for 1D parameters outperforms AdamW, future work could explore a hybrid approach—similar to Muon—where AdamW is used specifically for 1D parameters while maintaining ASGO for higher-dimensional structures. (ii) Employ exponential moving averages to update the preconditioner and incorporate momentum, following established practice, which results in Adam’s improvements across various tasks over AdaGrad. (iii) Compute and store the inverse of preconditioner matrix every $\tau$ iterations, as is done in distributed Shampoo (Shi et al., 2023), SOAP (Vyas et al., 2024) and Galore (Zhao et al., 2024). This modification increases memory requirements but substantially reduces computational overhead while improving algorithmic stability.

### v2

@@ 5 Smooth Theory
It is also important to study the performance of Algorithm 1 in smooth settings, as many real training tasks have been widely observed to be smooth, at least locally. Also, only in smooth settings can we properly describe the importance of batch size.
@@
If $X=[x_{1},\dots x_{n}]$, where each $x_{i}\in\mathbb{R}^{m}$, and we vectorize $X$, i.e., $\mathbf{x}\equiv\mathrm{vec}(X)=(x_{1}^{T},\ldots,x_{n}^{T})^{T}$, and form the block diagonal matrix $\mathbf{L\equiv}\mathrm{diag}[L,\dots,L]$, we obtain that ${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\mathbf{x}^{T}\mathbf{L}\mathbf{x}.$ This means that Assumption 2 is equivalent to the existence of a symmetric matrix $L\in\mathbb{R}^{m\times m},L\succ 0$ such that for any $w\in\mathbb{R}^{mn}$, $-\mathbf{L}\preceq\nabla^{2}f_{v}(\mathbf{w})\preceq\mathbf{L},$ where $f_{v}(\mathbf{w})=f(W)$, $W=[w_{1},\dots,w_{n}]\in\mathbb{R}^{m\times n}$ and $\mathbf{w}\equiv\mathrm{vec}(W)=[w_{1}^{T},\dots,w_{n}^{T}]^{T}\in\mathbb{R}^{mn}$. Hence, it is closely related to the block-wise diagonal structure of the Hessian as observed by many researchers; (see Figure 3 in Appendix A). Also note that Assumption 2 implies the standard smoothness assumption with respect to the Frobenius norm by $\left\|\nabla f_{v}(\mathbf{w})\right\|_{\mathrm{op}}\leq\left\|L\right\|_{\mathrm{op}}$. This block-wise diagonal smoothness is an extension of the standard smoothness, which is related to but different from the diagonal anisotropic smoothness employed for analyzing sign-based and adaptive gradient methods (Bernstein et al., 2018; Liu et al., 2024).
@@
Let $N_{t}\triangleq\nabla f(W_{t};\xi)-\nabla f(W_{t})\in\mathbb{R}^{m\times n}$ be the stochastic gradient noise. We assume that $\mathbb{E}[N_{t}]=0$ and there exists a symmetric positive definite matrix $V$ such that
$\mathbb{E}\left[N_{t}N_{t}^{\top}\right]\preceq V^{2}.$
One can check that Assumption 3 implies the standard variance bound $\mathbb{E}[\left\|N_{t}\right\|_{\mathrm{F}}^{2}]\leq\left\|V\right\|_{\mathrm{F}}^{2}$. The assumption shares some similarity with the coordinate-wise variance bounds in Bernstein et al. (2018); Crawshaw et al. (2022); Liu et al. (2024), in the sense that it allows a more fine-grained analysis. This matrix-form variance upper bound may better describe the real case since it takes the structure of the noise into account, which is relevant to matrix rank and other structured properties.
@@
Under Assumptions 1, 2 and 3, for Algorithm 1 with $\eta_{t}\equiv\eta=D_{\mathrm{op}}$ and a batch size of $M$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\leq\frac{4D_{\mathrm{op}}^{2}\left\|L\right\|_{*}}{T}+\frac{2\sqrt{2}D_{\mathrm{op}}\left\|V\right\|_{*}}{\sqrt{MT}}+\frac{2\epsilon D_{\mathrm{F}}^{2}}{D_{\mathrm{op}}T},$
where $D_{\mathrm{op}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{op}}$, and $D_{\mathrm{F}}\triangleq\max_{0\leq t\leq T-1}\left\|W_{t}-W_{*}\right\|_{\mathrm{F}}$.
As discussed in Section 4, $D_{\mathrm{op}}$ and $D_{\mathrm{F}}$ can be bounded if we add a projection step in the update. The convergence rate specified in Theorem 3 is similar to the rate in Corollary 2 if the batch size $M$ is small, when the $\mathcal{O}(1/\sqrt{MT})$ term dominates the rate, and thus shares the same properties as we discussed in Section 4. This means that ASGO can benefit if the stochastic gradient noise $V$ is generally low-rank. Furthermore, when the batch size $M$ is large so that the $\mathcal{O}(\left\|L\right\|_{*}/T)$ term contributes significantly to the bound, Theorem 3 implies more, which we now discuss.
Comparison with SGD. The convergence rate of SGD (Garrigos and Gower, 2023) is:
@@
where $D_{\mathrm{F}}$ is defined in Theorem 3. Since the comparison between the $\mathcal{O}(1/\sqrt{MT})$ term is generally consistent with the discussion in Section 4, we further compare the $\mathcal{O}(1/T)$ term here to provide some more intuition: we have
  • $\left\|L\right\|_{\mathrm{op}}\leq\left\|L\right\|_{*}\leq r_{L}\left\|L\right\|_{\mathrm{op}}$, where $r_{L}$ is the rank of $L$. Thus when $L$ is low-rank, or has very imbalanced singular values, $\left\|L\right\|_{*}$ can be close to $\left\|L\right\|_{\mathrm{op}}$;
  • $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}=$ max rank of $(W_{t}-W_{*})$. Thus when $W_{t}-W_{*}$ are of relatively high-rank or have lots of singular values of a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
Therefore, we can see that in general, ASGO should work well when the Hessian can be well approximated by a block-wise diagonal matrix with a low-rank $L$ for each block and $W_{t}-W_{*}$ are of relatively high-rank. Note that Hessians have been found to be low-rank in DNNs, especially after some steps of training (Sagun et al., 2016, 2017; Wu et al., 2020). Block-wise diagonal structure of DNNs. It has been widely observed that the Hessian of MLPs and other DNNs, are approximately block-wise diagonal (Collobert, 2004; Zhang et al., 2024a, b; Bahamou et al., 2022), as shown in Figure 3 in Appendix A. This block-wise diagonal structure naturally arises from the structure of MLPs, where all the parameters associated with a particular neuron (i.e., all elements of either a row or column of $W$) are more closely related and have a denser block Hessian matrix, than a set of parameters that do not share such an association. Intuitively, the elements of the rows of $W$ seem to be more likely to be closely related to the columns of $W$. Based on Theorem 3, our algorithm should perform well in this block Hessian setting, showing great potential in real applications. Intuition on single-side preconditioners. Since ASGO only uses single-side preconditioners, it is important to determine on which side they should be applied. As demonstrated in Figure 3, a block in the Hessian commonly corresponds to all input weights into a neuron in the next layer, i.e., as a row of $W$. Therefore, we may want the single-side preconditioner to be on the right side of $G_{t}$. Based on this block-wise diagonal structure, Shampoo-like double-sided preconditioners may not only suffer from higher memory and computation costs, but also may precondition less effectively, since they are trying to approximate the structured curvature information matrix using a dense matrix produced by a Kronecker product (Gupta et al., 2018; Morwani et al., 2024).
@@
The proof of Theorem 3 is given in Appendix F. The analysis is similar in its general outline to the smooth analysis for AdaGrad (Levy et al., 2018; Liu et al., 2024), but it is more complex because of the involvement of matrix operations in ASGO.
6 Further Discussions on ASGO
@@
Muon (Jordan et al., 2024) can be interpreted as a standard steepest descent algorithm utilizing the spectral norm (Bernstein and Newhouse, 2024b) with momentum. If we ignore the incorporated momentum, Muon computes
@@
preserving the matrix structure of the problem and naturally exploiting the structured properties because of the involvement of spectral norm in the steepest descent framework. This aligns with ASGO’s exploitation of structured properties. Moreover, if we ignore the momentum in both the gradient and the preconditioner, we can see that ASGO is equivalent to Muon. [^1] A proof of this equivalence can be found in Appendix G. Interestingly, this equivalence between ASGO and Muon in both theoretical basis and algorithm design is analogous to that between diagonal AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). In this sense, we may interpret ASGO and Muon as AdaGrad and SignSGD for structured gradients, respectively. Note that Shampoo also admits such a relation with Muon in algorithmic design. However, as discussed in Section 4, Shampoo has a worse convergence rate and thus fails to benefit from the structured properties as does ASGO. To some extent, this means that ASGO may be a more appropriate approach than Shampoo as such an analog to diagonal AdaGrad.
From this intuition, we may also conjecture what the nonconvex convergence rate of ASGO is based on what it is for Muon. Hence, we prove in Appendix G that this rate for Muon is:
$\min_{0\leq t\leq T-1}\left\|\nabla f(W_{t})\right\|_{*}^{2}\leq\mathcal{O}\left(\frac{\left\|L\right\|_{*}(f(W_{0})-f^{*})}{T}\right),$ (3)
where we assume that $f^{*}\triangleq\inf f(W)>-\infty$. Since diagonal AdaGrad and SignSGD have the same convergence rate in nonconvex settings up to logarithmic factors (Bernstein et al., 2018; Sun et al., 2023; Liu et al., 2024), we expect that ASGO has a nonconvex convergence rate comparable to (3) obtained by Muon up to logarithmic factors. It is an interesting future topic to prove this conjecture and further explore the nonconvex behavior of ASGO theoretically. Also, we note that as an intuitively smoother version of Muon, ASGO has good convergence properties in nonsmooth settings as shown in Section 4, where Muon, like SignSGD (Bernstein et al., 2018), may fail to converge (Karimireddy et al., 2019).
@@
However, in smooth settings, nonconvex convergence analysis of Muon has been presented in Li and Hong (2025). They establish the convergence results under a more general setting, involving gradient noise and momentum following the analysis in Cutkosky and Mehta (2020). However, if we only look at the deterministic case, their result is worse than (3) because of the explicit dependence on the dimension $n$. The key here is that the standard smoothness condition with respect to the Frobenius norm is not a good fit for analyzing structured gradient algorithms like Muon or ASGO. Using Assumption 2, we obtain better convergence results for Muon in (3).
Practical Implementations of ASGO.
We also provide Algorithm 2, a practical implementation of ASGO inspired by Adam (Kingma and Ba, 2014) and distributed Shampoo (Shi et al., 2023), together with a memory efficient diagonal variant of ASGO, named DASGO in Algorithm 3. DASGO is basically a light-weight optimizer with the preconditioner $\Lambda_{t}$ of ASGO being diagonalized, which is efficient in both memory and computations. Please refer to Appendix B for details.
7 Empirical Results
We empirically evaluate the effectiveness of ASGO (Algorithm 2) and DASGO (Algorithm 3) on pretraining and finetuning tasks for Large Language (LL) models. We compare the methods against established optimizers, including AdamW (Kingma and Ba, 2014; Loshchilov and Hutter, 2017), Shampoo (Gupta et al., 2018), and Muon (Jordan et al., 2024). Several important implementational details should be noted for fair comparison. First, since Muon is designed to operate exclusively on matrix parameters, we follow Jordan et al. (2024) and apply AdamW update rules to all 1D parameters within the Muon optimizer to ensure that it can handle the complete model. All experiments were conducted using NVIDIA V100s SMX2 GPUs. Specifically, for the larger-scale pretraining of GPT2, we utilized a configuration of four V100 GPUs, while other experiments were performed on a single V100 GPU.
7.1 Pretraining NanoGPT
We first conducted experiments using the NanoGPT architecture on the Shakespeare character dataset, following the configuration in (Karpathy, 2022). To ensure a fair comparison, we maintained consistent hyperparameter (HP) settings varying only those parameters that significantly impacted optimizer performance, i.e., learning rate, 1st and 2nd order moment coefficients ($\beta_{1}$ and $\beta_{2}$) and so on)  [^2] A full description of how these “best” HPs were selected and their values are presented in Appendix C.. For consistency across all optimization methods, we employ the OneCycleLR learning rate schedule, which has been shown to provide stable convergence properties in deep learning tasks. Figure 1 presents a comparison of the pretraining performance of ASGO and DASGO with that of Shampoo, Muon and AdamW on the NanoGPT model. See Appendix C for a plot of the training loss.
Figure 1: Validation Loss for NanoGPT
Examining both test and training loss curves, ASGO consistently outperforms Shampoo despite requiring only half the memory consumption and computational effort, which highlights ASGO’s practical advantages for training language models. Furthermore, ASGO and Muon achieve the lowest final training and test losses, outperforming all other methods. This consistent performance between ASGO and Muon aligns well with our discussions in Section 6. Moreover, The lightweight DASGO optimizer achieves competitive results against AdamW in both training and test loss metrics. DASGO demonstrates particularly strong performance during the initial training phase (first 200 training steps, which is about the first two epochs), but eventually exhibits a performance gap with ASGO. We will discuss this performance gap in Section 7.2.
7.2 Pretraining GPT2
To further evaluate the efficacy of ASGO and DASGO, we extended our investigation to larger-scale pretraining tasks. We adopted the configuration of the GPT2 model as described by (Karpathy, 2022). In this setting, the GPT2 model comprises 12 Transformer blocks, each with a hidden dimension of 768. Consequently, there are 12 projection layers with dimensions 768×2304. In each of these layers alone, Shampoo necessitates storing two matrices of size 768×768 and two matrices of size 2304×2304 for its state. In contrast, ASGO requires storing only two 768×768 matrices, corresponding to the preconditioning and inverse preconditioning factors. This substantial difference in memory requirements rendered Shampoo impractical for our large-scale GPT2 pretraining experiments, preventing us from obtaining meaningful results within our computational constraints. Therefore, we do not report Shampoo’s performance in this setting. To ensure a fair comparison, we carefully tuned the learning rates and $\beta_{2}$ values (where applicable) for all optimizers (tuning details are provided in Section C). Furthermore, we employed a learning rate schedule consisting of 200 linear warm-up steps followed by a cosine decay, which is a common practice for training LLM models. Figure 2a and Table 1 present a comparison of the training and validation loss achieved by ASGO and DASGO against Muon and AdamW on the GPT2 model.
Table 1: Pretraining GPT2 Validation Loss
AdamW Muon ASGO DASGO
3.96 3.82 3.87 5.23
The training loss curves reveal that ASGO achieves a lower final training loss compared to AdamW in the pretraining task, suggesting superior optimization. This observation is further corroborated by the validation loss (Table 1), where ASGO’s performance (3.87) is better than AdamW’s (3.96). Interestingly, ASGO’s performance appears more similar to Muon (3.82) in terms of final loss, although ASGO exhibits a slightly higher validation loss than Muon.
(a) Train Loss
(b) Learning Rate Sensitivity
Figure 2: Pretraining Performance and Learning Rate Sensitivity on GPT2
ASGO demonstrates potential advantages over Muon. As depicted in Figure 2a, ASGO shows a faster initial decrease in training loss. Furthermore, our learning rate sensitivity analysis (Figure 2b), with tested candidates $[10^{-3},5\times 10^{-3},10^{-2},3\times 10^{-2},5\times 1e^{-2},7\times 10^{-2},10^{-1},5\times 10^{-1}]$ using optimal settings for Muon and ASGO suggests that ASGO might be more robust to learning rate selection. Specifically, at higher learning rates (e.g., $10^{-1}$, $5\times 10^{-1}$), ASGO maintains relatively stable training while Muon exhibits instability. These observations – faster initial convergence and a potentially wider stable learning rate regime for ASGO – are consistent with characteristics often associated with adaptive optimization methods, as discussed in Section 6. However, DASGO (5.23) underperformed compared to AdamW, Muon, and ASGO in the GPT2 pretraining task. This performance gap was also observed in the pretraining of the smaller NanoGPT model. A primary reason for the difference between DASGO and ASGO likely stems from its diagonal preconditioning. By only retaining and using only the diagonal elements of the $GG^{T}$, DASGO essentially disregards the inter-dependencies between different parameter gradients from neurons within a layer, moving away from a true matrix-based adaptive approach towards a per-parameter scaling akin to vector-based methods. This highlights the value of preserving non-diagonal elements in the preconditioner matrices, particularly for capturing parameter interactions in attention-based architectures. Nevertheless, DASGO’s significantly reduced memory footprint makes it a compelling option for resource-limited settings where computational efficiency is paramount.
7.3 Finetuning GPT2-Large on WikiText-2
To further assess the performance of our proposed optimizers, we conducted fine-tuning experiments on the GPT2-Large(774M) model (Radford et al., 2019) using the WikiText-2 dataset. We explored two distinct fine-tuning objectives:
  • First, we fine-tuned GPT2-Large using the standard Causal Language Modeling (CLM) loss, which is the conventional approach for autoregressive language models.
  • Second, we also fine-tuned the same GPT2-Large model on WikiText-2 but employed the Fill-in-the-Middle (FIM) training objective, following the setting in (Bavarian et al., 2022). The FIM objective modifies the training process to enable the model to learn to infill text by rearranging document spans.
To account for statistical variability, we conducted five experimental runs with different random seeds under the same hyperparameter settings. Table 2 presents the average perplexity results after fine-tuning for 2 epochs. For both objectives, we employed a cosine decay learning rate scheduler. Hyperparameter search details and 95% confidence intervals are provided in Table 6.
Table 2: Finetuning GPT2-Large Perplexity
AdamW Muon ASGO DASGO
CLM 14.01 13.91 13.88 13.84
FIM 17.46 15.93 15.66 15.45
Under the CLM objective, ASGO (13.88 perplexity) and DASGO (13.84 perplexity) achieved lower perplexity than both Muon (13.91 perplexity) and AdamW (14.01 perplexity). This suggests that both ASGO and its memory-efficient variant, DASGO, can be effective for traditional LL modeling fine-tuning. When the model was fine-tuned using the FIM objective, we observed a similar trend, with ASGO (15.66 perplexity) and DASGO (15.45 perplexity) outperforming both Muon (15.93 perplexity) and AdamW (17.46 perplexity). Across both fine-tuning scenarios, ASGO and DASGO demonstrated competitive or superior performance compared to these baseline optimizers.
8 Conclusions
In this paper, we proposed a novel algorithm ASGO, which achieves significantly better convergence rates compared to full-matrix AdaGrad and Shampoo. Based on the theory, we demonstrated that ASGO can benefit from low-rank gradients and block-wise diagonal Hessians, which are widely observed structured properties of DNNs. We further proposed some practical modifications to ASGO, and verified its effectiveness empirically. Currently, ASGO still has two major limitations: 1) computationally intensive compared to Adam because of the matrix operation; 2) it is not straightforward to extend the algorithm to apply to tensors. We plan to look into these issues in future work.
Appendix A Additional Related Work
Optimization with Matrix Structure.
Much recent work has focused on improving full-matrix AdaGrad and Shampoo. Feinberg et al. [2023] uses a sketching-based approach to approximate the full-matrix AdaGrad preconditioner with lower memory cost. Morwani et al. [2024] provides theoretical intuition and empirical evidence to claim that Shampoo should use the $-1/2$ power in its preconditioners to better approximate the Empirical Fisher (EF) matrix. Vyas et al. [2024] demonstrates that Shampoo is like doing Adafactor in the eigenspace of gradients and proposes a novel algorithm, SOAP, that performs Adam in this eigenspace. SOAP is observed to achieve better performance than Adam and Shampoo, but suffers from a high computation load per iteration. Galore [Zhao et al., 2024] shares a similar algorithmic design with SOAP with extra focus on lowering memory costs. Muon [Jordan et al., 2024] follows this line of work, using a steepest descent (in the spectral norm) framework, which it shows to be scalable and effective in training large foundation models [Liu et al., 2025a]. Large et al. [2024], Bernstein and Newhouse [2024a] propose the modular approach that has also been applied to improve Muon. More recently, Nguyen et al. [2025] proposes AdaDiag, which may be viewed as SOAP doing SVD without gradient accumulation. Liu et al. [2025b] proposes COSMOS, a combination of SOAP and Muon, that trades off between performance and computational efficiency.
Rank of Gradients and Weight Updates.
It has been widely observed that gradients are naturally low-rank in DNNs, even when a large batch size is employed [Gur-Ari et al., 2018, Zhao et al., 2021, Yang et al., 2023, Cosson et al., 2023]. This property has been widely utilized for computation and memory efficiency in training [Wang et al., 2018, Cosson et al., 2023, Zhao et al., 2024]. On the other hand, the rank of the total weight update $\Delta W=W_{T}-W_{0}$, depends a lot on the training method. If we use LoRA [Hu et al., 2022], $\Delta W$ is determined to be low-rank. However, in pretraining or even many complex fine-tuning tasks, LoRA’s performance is much worse than methods that produce high-rank weight updates like full-parameter training, which is conjectured to be due to the weight update rank [Lialin et al., 2023, Jiang et al., 2024b, Huang et al., 2025].